this exam.
SHORT EXERCISE #1 (15 POINTS):
Tom’s Surf Company, whose fiscal year ends December 31, completed the following transactions involving notes payable:
2018
Nov 30 Purchased inventory display equipment by issuing a 60 day 10 percent note for $35,000. Dec 31 Made the end-of-year adjusting entry to accrue interest expense
2019
Jan 29 Paid off the balance of the note
Prepare the journal entries for the above transactions. Round your answers for interest calculations to the nearest cent. Assume there are 365 days in the year.
1
Expert's answer
2020-03-12T02:45:51-0400

Debit equipment credit current account, cash desk - $35,000 - equipment purchased

Debit other expenses credit loan calculations for loans: 574.58

"IC=C\times i\frac{Dm}{365}" = 35 000*0.1*31/365=297.26 -january 2018

"IC=C\times i\frac{Dm}{366}" = 35 000*0.1*29/366=277.32 - february 2019

IC - interest charges

C -credit

i - annual interest rate

Dm -number of days in a month

297.26+278.08=574.58

Debit loan calculations for loans credit balance of the note - 574.58